A gas mixture of 6 L of propaneand butane on complete combustion at 25 Celsius produce 20 L co2 Volcker of butane present initially in mixture is

Grade:11

1 Answers

Vikas TU
7 years ago
Since process has took as complete combustion concepts of POAC can be applied here.
No. of total carbons in Reactant = > 7
and let V be the volume of butane present in the initial mixture.
Therefore, Volume of propane = > 6 – V
Hence,
V * 4 + (6- V)*3 =  1 * 20
Solving for V we get,
4V – 3V = 20 – 18
V = 2 litre.
Hence butane is 2 litre in capacity.